OpenStudy (anonymous):

Evaluate the limit that exist limx->0 sin^2 3x/5x^2

OpenStudy (anonymous):

You have 0/0. Time to use l'hopital's.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \sin ^{2} 3x/ 5x ^{2}\]

OpenStudy (anonymous):

this the trigonomoetric limit that sinx/x = 1 and x/sinx = 1

OpenStudy (anonymous):

I'm not sure I'd trust that the limit of sin(3x)/x = limit of sin(x)/x, but it may be right.

OpenStudy (anonymous):

Nope, it's not.

myininaya (myininaya):

\[\frac{\sin3x*\sin3x}{5*x*x}=\frac{1}{5}*9*\frac{\sin3x}{3x}\frac{\sin3x}{3x}\]

myininaya (myininaya):

sin3x/3x->1 as x->0 1/5*9*1*1

myininaya (myininaya):

=9/5

OpenStudy (anonymous):

I still like l'hopital's better ;p

myininaya (myininaya):

lol

OpenStudy (anonymous):

and if you multiply bottom by 3 you have to multiply top by 3?

OpenStudy (anonymous):

To keep the same ratio, yes. That's where the 9 came from.

OpenStudy (anonymous):

wouldnt it just be 3/5?

OpenStudy (anonymous):

No because you need one 3 for each of the two threes you have in the denominator.

OpenStudy (anonymous):

so were multiplying by 9 for numerator and denominator then? ok thank you so much myininaya and polpak!

myininaya (myininaya):

np